OpenStudy (lifeisadangerousgame):

Could someone help me understand how to do this? Convert 3 ft. per minute into miles per hour. Show all work. Make sure you label your answer.

OpenStudy (shubhamsrg):

whats the relation between feet and mile?

OpenStudy (shubhamsrg):

if you can tell..

OpenStudy (lifeisadangerousgame):

There are 5,280 feet in a mile, and there are 60 seconds per minute

OpenStudy (lifeisadangerousgame):

in a minute I mean

OpenStudy (shubhamsrg):

if 1 mile = 5280 feet,,then 3 miles = 3*5280 feet! and 1 hr = 60 mins,,so 1 min = 1/60 hrs.. so,,can you reach the ans now?

OpenStudy (lifeisadangerousgame):

No sorry, I don't know what to do next. This is a new topic for me..

OpenStudy (anonymous):

3 feet per minute. 1 feet = 12 inches so 3 x 12 = 36 inches per minute. 1 inch = 2.54 cm so 36 x 2.54 = 91.44 cm per minute. 100 cm = 1meter so 91.44 / 100 = 0.9144 meter per minute. 1 hour = 60 minutes so if 0.9144 m in one minute..then in 60 minutes it is 60 x 0.9144=54.864 meter in one hour Now 1 mile = 1609 m so 54.864 meter = 0.034 mile so the 3 ft. per minute = 0.034 mile per hour

OpenStudy (shubhamsrg):

3 miles / 1 hr = 3*5280 feet / 60 mins solve 3*5280 /60 and you get your ans

OpenStudy (lifeisadangerousgame):

Ok, thank you!

OpenStudy (lifeisadangerousgame):

264?

OpenStudy (shubhamsrg):

yep..seems right..

OpenStudy (lifeisadangerousgame):

Ok, thank you!

OpenStudy (shubhamsrg):

sorry i stated it wrong above..really sorry @vikash4exploring did it correctly.. i quoted the ques wrong above,,really sorry.. but the method dosent change..

OpenStudy (anonymous):

shubha is wrong..

OpenStudy (lifeisadangerousgame):

oh, ok, thanks though, both of you :)

OpenStudy (lifeisadangerousgame):

Vikash, can you show me how you got it? I know you have everything there but like I said earlier, this is a new subject for me..

OpenStudy (anonymous):

i have already explained in detail above...what else do u need

OpenStudy (lifeisadangerousgame):

Mmk sorry, thanks :)

OpenStudy (anonymous):

If you need any further explanation..feel free to tell me

OpenStudy (lifeisadangerousgame):

Will do! Thanks again